ㄐ 3f X) y are ra ndon varial les) then Ahour that.
Answer #1

Cov( X+Y , X-Y) = Cov(X,X) - Cov(X,Y) + Cov( Y, X) - Cov(Y, Y)

= Var(X) - Var(Y)

Since Cov( X, X) = E(XX) -- E(X) E(X) = E(X​​​​​​2) - [ E(X) ]2 = Var(X)

Similarly Cov( Y, Y) = Var( Y)